OpenStudy (anonymous):

HELP!! SIMPLIFY.. √300x^16 Anwsers: A..10x^8√3 B..100x^8√3 C..150x^8 D..25x^4√12

OpenStudy (campbell_st):

look for the squares \[\sqrt{300x^{16}} = \sqrt{100\times(x^8)^2 \times3}\] take ot the squares \[10x^8\sqrt{3}\]

OpenStudy (anonymous):

Thank u...haha is your last name campbell?

OpenStudy (campbell_st):

nope sorry

OpenStudy (anonymous):

ok